# A Bayesian framework for active object recognition, pose estimation and shape transfer learning through touch

### v1

@@ Abstract
As humans can explore and understand the world through the sense of touch, tactile sensing is also an important aspect of robotic perception. In unstructured environments, robots can encounter both known and novel objects, this calls for a method to address both known and novel objects. In this study, we combine a particle filter (PF) and Gaussian process implicit surface (GPIS) in a unified Bayesian framework. The framework can differentiate between known and novel objects, perform object recognition, estimate pose for known objects, and reconstruct shapes for unknown objects, in an active learning fashion. By grounding the selection of the GPIS prior with the maximum-likelihood-estimation (MLE) shape from the PF, the knowledge about known objects’ shapes can be transferred to learn novel shapes. An exploration procedure with global shape estimation is proposed to guide active data acquisition and conclude the exploration when sufficient information is obtained. The performance of the proposed Bayesian framework is evaluated through simulations on known and novel objects, initialized with random poses and is compared with a rapidly explore random tree (RRT).The results show that the proposed exploration procedure, utilizing global shape estimation, achieves faster exploration than the RRT-based local exploration procedure. Overall, results indicate that the proposed framework is effective and efficient in object recognition, pose estimation and shape reconstruction. Moreover, we show that a learned shape can be included as a new prior and used effectively for future object recognition and pose estimation of novel objects.
@@
Bayesian inference, robotics, tactile perception, active learning, object recognition, pose estimation, shape reconstruction
@@ I Introduction
Tactile sensing is an important aspect of robot sensing, and it is useful both to complement vision, but also as a standalone sensing modality. Visual information is not always available and reliable, especially under severe occlusions and poor lighting conditions. Therefore, methods that can work regardless of the visual information availability are desired, to have robots function well in natural or unstructured environments. In this study, we focus on localizing and recognizing objects based on the shape of the object, solely through touch. Knowing the object’s shape and pose is important for locating and manipulating objects, and therefore, it is useful for a variety of applications.
@@
In unstructured environments, both known and novel objects can be encountered, nonetheless, how to address both known and novel objects under an unified framework remains an open question. Identifying known and unknown objects, and addressing them accordingly is therefore of paramount importance. For known objects, the object class and pose should be estimated, whereas, for a novel object, its shape should be explored and learned. Previous work tends to separate the processes of object recognition, pose estimation, and shape reconstruction. In this study, we combine these goals. Moreover, we explore the possibility of transferring knowledge from known object shapes to learn novel shapes by addressing these novel objects within the same framework that has obtained prior knowledge about objects.
Compared to robotic vision, robotic tactile sensors provide sparse data that are often insufficient for resolving uncertainty in the object class/shape and pose estimation. Especially when the tactile sensors are small compared to the size of the object in contact, only local information is available. Therefore, tactile exploration needs to be performed in an active manner to reduce the uncertainty in object class, pose and shape.
With the recent advancement of tactile sensing technology, vision-based tactile sensors [1, 2] have gained popularity, as they provide accurate and detailed information on the local geometry of the contact area, although they are still less accessible due to their relatively high cost and complicated manufacturing procedure. In this study, we assume the input to the system to be a contact mask with the sensor location and a surface normal vector when in contact. The sensor’s location and orientation can be extracted from robot’s proprioception through forward kinematics, while the normal vector can be estimated based on the sensor’s orientation.
@@
A Bayesian-based approach is a good choice for the problem introduced in this study, because it allows for straightforward integration of prior knowledge with sensor data, and it has an explicit uncertainty treatment. Therefore, we use a Bayesian framework to jointly estimate the object class and object pose with active exploration from first contact to conclusion. Moreover, the proposed framework can identify and learn the shape of novel objects through Bayesian evidence tracking and transfer learning.
@@
An outline of the proposed framework is shown in Fig. 1. The particle filter (PF) is in charge of online updating the joint distribution of object class and pose, while more data are collected. Taking inspiration from the Manifold particle filter [3] and point-pair features proposed in [4], we modify a traditional particle filter [5] to progressively sample new particles based on newly observed data points. Therefore, our particle filter can deal with arbitrary initial uncertainty. In addition, this approach keeps the computational cost tractable at each time step and can achieve low pose estimation error as more data are collected despite using coarse discretized approximation in the PF. The Bayesian evidence of the maximum likelihood estimation (MLE) obtained from the PF is used to determine if the object is a known or novel object. Once a novel object is detected, the MLE from the PF is used as a prior for Gaussian process implicit surface (GPIS) reconstruction. The target point selection and contact enforcement modules take care of the active data acquisition using information from global shape estimation from either the PF or GPIS. A termination criterion is used, based on the Directed Hausdorff Distance(DHD), to determine when the object has been sufficiently explored and a conclusion can be drawn.
@@
Figure 1: The proposed framework consists of a particle filter (PF) and a Gaussian process implicit surface (GPIS). The maximum likelihood estimation (MLE) for object class and pose from the PF are used as priors for GPIS reconstruction when a novel object is identified. An exploration procedure, including target point selection and contact enforcement, is proposed to perform active data acquisition.
@@
  • A Bayesian framework is proposed to perform simultaneous object recognition and pose estimation for known objects. In case of novel objects shape transfer learning is applied in an active setting to learn the new shapes.
  • The proposed Bayesian framework can capitalize on the knowledge of a learned novel object shape for faster object recognition and pose estimation when the object is later encountered.
@@ II-A Bayesian Touch-Only Object Recognition and Pose Estimation
Previous work approached the object recognition problem based on material properties. Kaboli et al. and Xu et al.[6, 7, 8] used properties such as stiffness, texture, and thermal conductivity to distinguish between different objects. Our method instead focuses on object recognition with geometric information, which can be obtained through most tactile sensors.
Some researchers used methods more in line with our work with touch-based pose estimation of objects. For instance, particle filters (PF) can be used to approximate arbitrary distributions, which makes them suited for capturing the uncertainty of the object pose presented during tactile exploration. However, using a PF comes with a computational challenge as a substantial amount of particles is required to cover the state space. Koval et al. [3] proposed the manifold particle filter by adaptively sampling particles that reside on the contact manifold to increase sampling efficiency. Petrovskaya et al. [9] proposed the Scaling Series algorithm, combining the Bayesian Monte Carlo technique coupled with annealing, to refine the posterior distribution of the 6-degree-of-freedom (DOF) pose through multiple stages with a small number of particles at each stage. Vezzani et al. [10] proposed to use a memory unscented Kalman filter (MUKF) to cover the 6-DOF pose space with a small number of particles where each particle is treated as a Gaussian distribution instead of a discrete sample. Although the methods are quite efficient, the work by [3, 9, 10] was confined to pose estimation. As they limited their scope to the pose estimation of a known object, multi-class classification was not addressed in their work. Vezzani et al. [11] extended their own work on pose estimation [10] to also address object recognition by applying the localization scheme to multiple object classes and picking the object class with the smallest error in localization.
Differing from the aforementioned methods, our proposed framework estimates the joint distribution of object class and object pose by progressively sampling new particles based on newly observed point pairs. The motivation behind this design choice is to avoid tracking all possibilities from the beginning of the exploration. Instead of using all data, the particle filter samples new particles that fit a new point pair generated from the new contact point and a previous observed contact point. Due to the discretized nature of the PF, there is no guarantee that a small number of initial particles can cover the region near the ground truth. By sampling new particles at each time step, the PF can consider more particles throughout active tactile exploration while keeping the number of particles small at each time step through resampling. More importantly, we combine the PF and GPIS to distinguish between known and novel objects and transfer knowledge on known object shapes to learn novel object shapes. Another distinction is that our proposed framework uses an active exploration procedure utilizing the global estimated shape at each time step for active data acquisition.
@@ II-B Bayesian Touch-Only Shape Reconstruction
Gaussian processes (GPs) are widely used for the task of shape reconstruction as they provide uncertainty measurement of the reconstructed surface, which enables Bayesian optimization for selecting actions during exploration. Sun et al. and Jamali et al.[12, 13] used the Gaussian process explicit surface for shape reconstruction, but the choice of representation cannot represent arbitrary shapes. Gaussian process implicit surface (GPIS)[14] was introduced to address this limitation. Several researchers, e.g.[15, 16, 17, 18, 19, 20], adopted GPIS for active shape reconstruction based on touch. These methods used the same prior for all test objects, which did not explore the possibility of shape transfer learning. Martens et al. [21] proposed to fit a parametric ellipsoid according to the observed point cloud as prior for GPIS reconstruction, whereas our approach updates the maximum likelihood prior continuously to take on the shape of objects in the prior at each time step based on new observations.
@@ II-C2 Active Exploration Strategies for Shape Reconstruction
Several studies used GP-based approaches. Jamali et al.[13] and Yi et al. [23] proposed active learning procedures that select the next point of interest to be the point with the most uncertainty. Yang et al.[24] proposed a greedy target point selection procedure based on mutual information of the GPIS. Using graph-based planning, Matsubara et al. [25] took into account both surface uncertainty and travel cost to improve the efficiency of tactile exploration. Instead of considering the uncertainty of each point, Driess et al. [18] proposed to generate smooth trajectories with the most uncertainty on the object’s surface for more human-like tactile exploration. Their subsequent study [19] introduced a new loss function based on the differential entropy of the GP to enable multi-finger exploration simultaneously. One problem with GP-based continuous path exploration is that it requires many points to get the global context of the object. In our study, motivated by obtaining global context as soon as possible, the most uncertain point on the GPIS will be selected as the target point when a novel object is detected.
For both known and novel objects, a contact enforcement procedure is performed to attempt to establish contact from the point of interest on the estimated shape. It will resort to local exploration if it fails to establish contact based on the estimated shape.
@@ II-D Deep-Learning-Based Approaches for shape reconstruction
Deep learning methods were developed to address active shape reconstructions. Comi et al. [26] learned a compact vector representation from large object datasets, and created a DeepSDF network conditioned on the learned representation. Based on contact patches from a vision-based tactile sensor, they retrieved the reconstructed shape by finding the shape embedding that yields the most consistent results with the contact. Their work does not consider an active learning setup. Wang et al. [27] and Rustler et al. [28] developed methods to perform active shape completion based on touch data, but their methods relied on information from RGB/RGB-D camera for initialization. Smith et al. [29, 30] proposed graph-neural-network-based methods (GNNs) to reconstruct 3D shapes using vision and touch. Though aiming at visuo-tactile information integration, their methods also work under touch-only settings. In addition, Smith et al. [30] used reinforcement learning to train an active exploration policy for efficient shape reconstruction. However, in contrast to our work, they used a fixed spherical prior for shape reconstruction. The aforementioned methods did not provide the joint distribution of the object class and object pose nor addressed the uncertainty throughout tactile exploration explicitly. In addition, they require a significant amount of training data.
@@ III Methods
In this section, we start by deriving the Bayesian formulation of the object recognition and pose estimation problem, then expand the formulation to include shape reconstruction for novel objects.
III-A Problem Formulation
To estimate the object class and object pose through tactile contact points and surface normal vectors at these points, we consider the joint distribution of the object class and object pose, given the tactile observations., since both object class and pose determine whether, or not, contact with the object can be observed at a certain location. Often an object is held stationary by one hand and explored by the other hand, therefore, for simplicity, we assume that the object’s pose is stationary during tactile exploration of the robotic arm.
The object class $c$ and the object pose $\boldsymbol{p}\leftarrow[p_{x},p_{y},p_{z},\alpha,\beta,\gamma]$ are concatenated into a latent variable vector $\boldsymbol{z}$, where $[p_{x},p_{y},p_{z}]$ represents the origin coordinates of the object frame in the world frame and $[\alpha,\beta,\gamma]$ denote the relative yaw, pitch, and roll angles of the object frame in the world frame. Each observed data point coordinate $[u_{x},u_{y},u_{z}]^{T}$ is denoted as $\boldsymbol{x}$. Likewise, the observed signed distance value $d_{s}$ and surface normal vector $[n_{x},n_{y},n_{z}]$ at each data point $\boldsymbol{x}$, is concatenated into a vector $\boldsymbol{d}$, i.e.,
$\boldsymbol{z}\leftarrow\left[c,\boldsymbol{p}\right]$ (1)
$\boldsymbol{d}\leftarrow\left[d_{s},\boldsymbol{n}\right]$ (2)
$\begin{cases}d_{s}=0&\text{on surface}\\ d_{s}>0&\text{outside object}\\ d_{s}<0&\text{inside object }\end{cases}$ (3)
Given an object class $c$ and object pose $\boldsymbol{p}$, the predicted Euclidean signed distance value $\widehat{d}_{s}$ from a point $\boldsymbol{x}$ to the object surface is given by its signed distance function
$\widehat{d}_{s}:=f(\boldsymbol{x},\boldsymbol{z}),$ (4)
where $f$ denotes the Euclidean signed distance function.
The predicted surface normal $\boldsymbol{\widehat{n}}$ at a contact point $\boldsymbol{x}$ could be calculated by differentiation, i.e.,
$\boldsymbol{\widehat{n}}=\nabla_{\boldsymbol{x}}{f}(\boldsymbol{x},\boldsymbol{z}).$ (5)
The predicted value of $\boldsymbol{d}$ is denoted as $\boldsymbol{\widehat{d}}$, where
$\boldsymbol{\widehat{d}}\leftarrow\left[\widehat{d}_{s},\boldsymbol{\widehat{n}}\right].$ (6)
@@ III-A Problem Formulation
III-B Bayesian Inference for Simultaneous Object Recognition and Pose Estimation and Shape Reconstruction through Touch

### v3

# A Bayesian Framework for Active Tactile Object Recognition, Pose Estimation and Shape Transfer Learning

@@ Abstract
As humans can explore and understand the world through active touch, similar capability is desired for robots. In this paper, we address the problem of active tactile object recognition, pose estimation and shape transfer learning, where a customized particle filter (PF) and Gaussian process implicit surface (GPIS) is combined in a unified Bayesian framework. Upon new tactile input, the customized PF updates the joint distribution of the object class and object pose while tracking the novelty of the object. Once a novel object is identified, its shape will be reconstructed using GPIS. By grounding the prior of the GPIS with the maximum-a-posteriori (MAP) estimation from the PF, the knowledge about known shapes can be transferred to learn novel shapes. An exploration procedure based on global shape estimation is proposed to guide active data acquisition and terminate the exploration upon sufficient information. Through experiments in simulation, the proposed framework demonstrated its effectiveness and efficiency in estimating object class and pose for known objects and learning novel shapes. Furthermore, it can recognize previously learned shapes reliably.
@@
Bayesian inference, tactile perception, active learning, object recognition, pose estimation, shape reconstruction
@@ I Introduction
Tactile sensing is an important aspect of robot sensing, because it can be used both as a complementary sense for vision and as a standalone sensing modality. Visual information is not always available and reliable, especially under severe occlusions and poor lighting conditions. Therefore, methods that can work regardless of the visual information availability are desired to have robots function well in unstructured environments.
Compared to robotic vision, robotic tactile sensors provide sparse data that are often insufficient for resolving uncertainty in robotic perception due to the lack of global context. Therefore, tactile exploration needs to be performed in an active manner to acquire sufficient information. To make our framework applicable to most tactile sensors, the input to the system is assumed to be a contact mask with the sensor location and a surface normal vector when in contact. The sensor’s location and orientation can be extracted from robot’s proprioception through forward kinematics, while the normal vector can be estimated based on the sensor’s orientation if the sensor does not provide normal force measurement.
In this study, we focus on localizing and recognizing objects based on the shape of the object through active tactile exploration. Knowing the object’s shape and pose is important for locating and manipulating objects, and therefore, it is useful for a variety of applications.
@@
In unstructured environments, both known and novel objects can be encountered. Thus identifying known and novel objects, and addressing them accordingly is of paramount importance. For a known object, the object class and pose should be estimated, whereas, for a novel object, its shape should be explored and learned. Nonetheless, how to address both known and novel objects under an unified framework remains an open question. Previous work tends to address known objects and novel objects separately. In this study, we explore the possibility of transferring knowledge from known object shapes to learn novel shapes by addressing novel objects within the same framework that has obtained prior knowledge about known objects.
@@
A Bayesian-based approach is a good choice for the problem introduced in this study, because it allows for straightforward integration of prior knowledge with sensor data, and it has an explicit uncertainty treatment. To this end, we propose a Bayesian framework to jointly estimate the object class, object pose and shape with active exploration from first contact to conclusion. Specifically, the proposed framework can identify and learn the shape of novel objects through Bayesian model evidence tracking and transfer learning.
@@
An outline of the proposed framework is shown in Fig. 1. At each time step, a customized particle filter (PF) is used to update the joint distribution of object class and pose based on newly obtained data. The Bayesian evidence of the maximum a posteriori (MAP) combination of object class and object pose obtained from the PF is then used to determine whether the object is known or novel. Once a novel object is detected, the MAP from the PF is used as a prior to perform Gaussian process implicit surface (GPIS) reconstruction. Based on the global shape estimation from either the MAP or GPIS, the target point selection and contact enforcement modules carry out active data acquisition. A termination criterion based on the Directed Hausdorff Distance (DHD) is applied to determine when the object has been sufficiently explored for automatic termination.
@@
Figure 1: The proposed framework consists of a customized particle filter and a Gaussian process implicit surface (GPIS). The customized particle filter estimate the joint distribution of object pose and object class upon newly obtained tactile observation, from which the maximum a posteriori (MAP) combination of object class and pose can be extracted. The MAP is used as a prior for GPIS reconstruction when a novel object is identified. An exploration procedure based on the global shape estimation (MAP/GPIS), including target point selection and contact enforcement is proposed to perform active data acquisition. Tactile exploration continues until the termination criterion is met. The learned GPIS from a novel object can be added as a new prior, which enables the framework to recognize it in future exploration.
@@
  • We proposed a unified Bayesian framework for active tactile object recognition, pose estimation and shape reconstruction.
  • The proposed Bayesian framework can discriminate between known and novel objects, estimate class and pose for known objects, while transferring knowledge on known shapes to learn novel shapes.
  • The proposed Bayesian framework can extend its prior knowledge with the learned shape and recognize this shape in future exploration.
@@ II-A Bayesian Touch-Only Object Recognition and Pose Estimation
Some previous work approached the object recognition problem based on material properties. Kaboli et al. and Xu et al.[1, 2, 3] used properties such as stiffness, texture, and thermal conductivity to distinguish different objects. Our method instead focuses on object recognition with geometric information, which can be obtained through most tactile sensors.
Methods in line with our work are associated with touch-based pose estimation of objects. Koval et al. [4] proposed the manifold particle filter by adaptively sampling particles that reside on the contact manifold to increase sampling efficiency. Petrovskaya et al. [5] proposed the Scaling Series algorithm, combining the Bayesian Monte Carlo technique coupled with annealing, to refine the posterior distribution of the 6-degree-of-freedom (DOF) pose through multiple stages with a small number of particles at each stage. Vezzani et al. [6] proposed to use a memory unscented Kalman filter (MUKF) to cover the 6-DOF pose space with a small number of particles where each particle is treated as a Gaussian distribution instead of a discrete sample. Although these methods are quite efficient, they were confined to the pose estimation problem. As they limited their scope to the pose estimation of a known object, multi-class classification was not considered. Vezzani et al. [7] extended their own work on pose estimation [6] to address object recognition by applying the localization scheme to multiple object classes and picking the object class with the smallest error in localization.
Differing from the aforementioned methods, our proposed framework estimates the joint distribution of object class and object pose, which is computationally challenging for a traditional particle filter[8]. Taking inspiration from the Manifold particle filter [4] and point-pair features proposed in [9], we propose a customized particle filter to progressively sample new particles based on newly observed data point pairs. This approach keeps the number of particles tractable at each time step and achieves the desired pose estimation error as more data are collected despite using coarse discretized approximation in the PF. More importantly, we combine the PF and GPIS to discriminate between known and novel objects and transfer knowledge on known object shapes to learn novel object shapes. Another distinction is that our proposed framework uses an active exploration procedure utilizing the global estimated shape at each time step for active data acquisition.
@@ II-B Bayesian Touch-Only Shape Reconstruction
Gaussian processes (GPs) are widely used for the task of shape reconstruction as they provide uncertainty measurement of the reconstructed surface, which enables Bayesian optimization for selecting actions during exploration. Jamali et al. [10] used the Gaussian process explicit surface for shape reconstruction, but their choice of representation cannot represent arbitrary shapes. Therefore, the Gaussian process implicit surface (GPIS)[11] was used to address this limitation. Several researchers, e.g.[12, 13, 14, 15, 16, 17], adopted GPIS for active shape reconstruction based on touch. These methods used the same prior for all test objects, which did not explore the possibility of shape transfer learning. Martens et al. [18] proposed to fit a parametric ellipsoid according to the observed point cloud as prior for GPIS reconstruction, whereas our approach updates the maximum likelihood prior continuously to take on the shape of known objects at each time step based on new observations.
@@ II-C2 Active Exploration Strategies for Shape Reconstruction
Several studies used GP-based approaches. Jamali et al.[10] and Yi et al. [20] proposed active learning procedures that select the next point of interest to be the point with the most uncertainty. Yang et al.[21] proposed greedy target point selection procedure based on mutual information of the GPIS. Using graph-based planning, Matsubara et al. [22] took into account both surface uncertainty and travel cost to improve the efficiency of tactile exploration. Instead of considering the uncertainty of each point, Driess et al. [15] proposed to generate smooth trajectories with the most uncertainty on the object’s surface for more human-like tactile exploration. Their subsequent study [16] introduced a new loss function based on the differential entropy of the GP to enable multi-finger exploration simultaneously. In our study, similar to the work from Jamali et al.[10] and Yi et al. [20], the most uncertain point on the GPIS will be selected as the target point to obtain global context as soon as possible.
Once a target point is determined, a contact enforcement procedure is performed to establish contact from the point of interest on the estimated shape. When failing to establish contact, the framework will record non-contact points and resort to a local exploration procedure.
@@ II-D Deep-Learning-Based Approaches for shape reconstruction
Deep-learning-based methods were developed to address active shape reconstructions. Comi et al. [23] learned a compact vector representation from large object datasets, and created a DeepSDF network conditioned on the learned representation. Based on contact patches from a vision-based tactile sensor, they retrieved the reconstructed shape by finding the shape embedding that yields the most consistent results with the contact. Their work does not consider an active learning setup. Wang et al. [24] and Rustler et al. [25] developed methods to perform active shape completion based on touch data, but their methods relied on information from RGB/RGB-D camera for initialization. Smith et al. [26, 27] proposed graph-neural-network-based methods (GNNs) to reconstruct 3D shapes using vision and touch. Though aiming at visuo-tactile information integration, their methods also work under touch-only settings. In addition, Smith et al. [27] used reinforcement learning to train an active exploration policy for efficient shape reconstruction. However, in contrast to our work, they used a fixed spherical prior for shape reconstruction. The aforementioned methods did not provide the joint distribution of the object class and object pose nor addressed the uncertainty throughout tactile exploration explicitly. In addition, they require a significant amount of training data.
@@ III Methods
In this section, we start by deriving the Bayesian formulation of the object recognition and pose estimation problem for known objects, then expand the formulation to enable shape transfer learning for novel objects. Both follow an exploration procedure for active data acquisition.
@@
III-A Bayesian Formulation of Object Recognition and Pose Estimation through Touch
In this study, we estimate the joint distribution of the object class and object pose, given the tactile observations, because it is the combination of the object class and object pose that gives rise to the observed tactile signal at a certain location. For simplicity, we assume that the object is rigid, and that the object’s pose is stationary during tactile exploration.
III-A1 Latent Variables
The object class denoted by $c$ is a categorical variable, while the object pose, denoted by $\mathbf{p}$, is a vector that contains the translation and the Euler angles of the object (yaw, pitch, and roll) with respect to the world frame. For compactness of the notations, the object class $c$ and the object pose vector $\mathbf{p}$ are concatenated into a latent variable vector $\mathbf{z}$, i.e.,
$\mathbf{z}\leftarrow\left[c,\mathbf{p}\right].$ (1)
III-A2 Object Priors
Our framework encodes the knowledge of known object shapes in the form of signed distance functions (SDFs) [28]. Given an object class $c$ and object pose $\mathbf{p}$, the predicted signed distance $\widehat{d}_{s}$ from a point $\mathbf{x}$ to the object surface can be obtained through the signed distance function
$\widehat{d}_{s}:=f(\mathbf{x},\mathbf{z}).$ (2)
$\begin{cases}\widehat{d}_{s}=0&\text{on surface}\\ \widehat{d}_{s}>0&\text{outside object}\\ \widehat{d}_{s}<0&\text{inside object }\end{cases}$ (3)
The predicted surface normal vector $\mathbf{\widehat{n}}$ at a contact point $\mathbf{x}$ could be calculated by differentiation, i.e.,
$\mathbf{\widehat{n}}=\nabla_{\mathbf{x}}{f}(\mathbf{x},\mathbf{z}).$ (4)
III-A3 Observations
In this study, each tactile sensor is abstracted as a single point. Each tactile data point consists of the sensor location $\mathbf{x}$ in the world frame, the signed distance value $d_{s}$ and surface normal vector $\mathbf{n}$ observed at location $\mathbf{x}$. The observed signed distance value $d_{s}$ can be inferred from the contact mask, i.e.,
$\begin{cases}d_{s}=0&\text{contact}\\ d_{s}>0&\text{no contact}\\ \end{cases}.$ (5)
For compactness, $d_{s}$ and $\mathbf{n}$ is concatenated into a observation vector $\mathbf{d}$, i.e.,
$\mathbf{d}\leftarrow\left[d_{s},\mathbf{n}\right]$ (6)
Similarly, the predicted value of $\mathbf{d}$ is denoted as $\mathbf{\widehat{d}}$, where
$\mathbf{\widehat{d}}\leftarrow\left[\widehat{d}_{s},\mathbf{\widehat{n}}\right].$ (7)
III-A4 Bayesian Inference of Object Class and Object Pose